OpenStudy (seattle12345):

Reduce cos2(6u) to a trigonometric function raised to the first power.

OpenStudy (anonymous):

Remember that: \[ \cos(2x)=\cos^2(x)-\sin^2(x) \]And, of course: \[ \sin^2(x)+\cos^2(x)=1 \]See how you can manipulate this to get a useful formula.

OpenStudy (seattle12345):

sorry it should be \[\cos ^{2}\left( 6u \right)\]

OpenStudy (seattle12345):

but maybe I'm blind because I don't see how to reduce it to the first power

OpenStudy (anonymous):

Let me help you out a bit: \[ \cos(2x)=\cos^2(x)-\sin^2(x)=\cos^2(x)-(1-\cos^2(x))=2\cos^2(x)-1 \]Use this to reduce the power.

OpenStudy (seattle12345):

would I put that into the power reducing formula?